Find the mean, variance, and standard deviation of n=90, p=0.2
Answer by stanbon(26259)

Find the mean, variance, and standard deviation of n=90, p=0.2
-------------------
mean = np = 90*0.2 = 18
variance = npq = 18*0.8 = 14.4
std = sqrt(npq) = 3.795
==============================
Cheers,
Stan H.